OpenStudy (anonymous):

4(z − 3) ≤ 3(z + 5) write as an interval notation... I got that x=27

jimthompson5910 (jim_thompson5910):

Keep in mind that there is a ≤ sign and not a = sign

jimthompson5910 (jim_thompson5910):

IF the equation was 4(z − 3) = 3(z + 5) then you would be correct in saying that the answer is x = 27

OpenStudy (anonymous):

but how do I write that in interval notation.

jimthompson5910 (jim_thompson5910):

You first have to figure out if it's x <= 27 or x >= 27

jimthompson5910 (jim_thompson5910):

So use the same techniques you used to solve the equation, but keep the inequality sign

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

did you mean negative infinity?

OpenStudy (anonymous):

YES...OH THANK YOU

OpenStudy (anonymous):

I have a few more like this that I sure could use coaching on. I am still shaky on the subject...so if you would be willing to hang with me here for a little bit for some support, I truly would appreciate it

jimthompson5910 (jim_thompson5910):

sure i can answer a few more

OpenStudy (anonymous):

\[2x-8>-2\] .... I came up with -3... would I write it as [-3.00)

jimthompson5910 (jim_thompson5910):

keep the inequality sign and flip it if you multiply/divide both sides by a negative number 2x - 8 > -2 2x > -2+8 2x > 6 x > 6/2 x > 3 So it should be (3, infinity) If there is a line under the inequality sign, then it is [3, infinity)

OpenStudy (anonymous):

Let me see if I understand this... if it is a definitive > without the hashtag under it... the answer is a ( whereas the > with the hashtag under it, it is a [

jimthompson5910 (jim_thompson5910):

that is correct, by hashtag you mean underline I've never heard of "hashtag" being used in this context, but underline sounds more appropriate here

OpenStudy (anonymous):

you are correct.. I just couldn't think of the word.. this math has me frazzled... That said... does this next problem make any sense to you... h = 48t + 1/2 at^2

jimthompson5910 (jim_thompson5910):

what do you want to do with it?

OpenStudy (anonymous):

solve the formula for a specified variable... for 'a'

jimthompson5910 (jim_thompson5910):

The idea is to isolate 'a' so we must first subtract 48t from both sides then multiply both sides by 2 finally, we divide both sides by t^2 to isolate 'a'

jimthompson5910 (jim_thompson5910):

\[\Large h = 48t + \frac{1}{2}at^2\] \[\Large h - 48t = \frac{1}{2}at^2\] \[\Large 2(h - 48t) = at^2\] \[\Large \frac{2(h - 48t)}{t^2} = a\] \[\Large a = \frac{2(h - 48t)}{t^2}\]

OpenStudy (anonymous):

HOLY GOD... what is THAT???

jimthompson5910 (jim_thompson5910):

the steps in solving for 'a'

OpenStudy (anonymous):

I understand that... but the mere sight of that is intimidating...